Refrigerant-134a is to be cooled by water in a condenser. The refrigerant enters the condenser with a mass flow rate of 6 kg/min at 1 MPa and 70ºC and leaves at 35°C. The cooling water enters at 300 kPa and 15°C and leaves at 25ºC. Neglecting any pressure drops, determine (a) the mass flow rate of the cooling water required and (b) the heat transfer rate from the refrigerant to water


1
Expert's answer
2022-02-14T00:01:02-0500

Part A


Taking the compressed water as a saturated liquid:

Internal energy at "15°C,\>\>h_{f1}=62.982kJkg^{-1}"


Internal energy of water at 25° C,

"h_{f2}" "=104.83kJ\/kg"


The refrigerant enters the condenser as superheated vapor at 70°C and leave as compressed liquid at 35°C.


Internal energy of superheated vapor R -134a at 70°C

"h_{f3}" "=303.87kJ\/kg"


Internal energy of compressed liquid R-134a at 35°C

"h_{f4}=100.88kJ\/kg"


Taking "m_w=" Flow rate of water


"m_w(104.83-62.982)=6(308.87-100.88)"


"m_w=29.1kg\/min"


Part B

Change in energy of water "=(104.83-62.982)kJ\/kg"

Rate of flow of water "=29.1kg\/min"


Heat transfer rate "=29.1\u00d7(104.83-62.982)"

"=1217.8kJ\/min"